OpenStudy (anonymous):

Bill has 2/3 of his pizza left from lunch. If he eats 3/7 more pizza, how much pizza is left?

OpenStudy (jdoe0001):

\(\bf \cfrac{2}{3}-\cfrac{3}{7}=\cfrac{}{{\color{brown}{ lcd?}}}\)

OpenStudy (anonymous):

can you explain it more

OpenStudy (anonymous):

Well, you need to make the denominators the same. The least common denominator would be 21. So mulitply 3 and 2 by 7, and muliply 3 and 7 by 3. So, we'd have: \[\frac{ 14 }{ 21 } - \frac{ 9 }{ 21 } = \frac{ 5 }{ 21 }\]

OpenStudy (anonymous):

5/21 is the answer. (:

OpenStudy (anonymous):

Do you understand now?

OpenStudy (anonymous):

yes thank you so much

OpenStudy (anonymous):

You're welcome!